Find the global maximum and minimum values of the following function on the given interval. If there are multiple points in a single category list the points in increasing order in x value and enter N in any blank that you don't need to use.

f(x)=4e−x−4e−2x  [0,1]



Global maxima

x = ____ Y=___

x=_____ Y=____

x=____ Y=___

Global minima

x=____ Y=___

x=____ Y=___

x=___ Y=___

x=____ Y=___






1
Expert's answer
2022-04-05T07:15:47-0400
"f(x)=4e^{−x}−4e^{−2x} \qquad[0,1]\\
\begin{aligned}
f'(x) =& -4e^{-x}+8e^{-2x}
\end{aligned}"


Find the critical points:


"\begin{aligned}
f'(x)=0 \implies -4e^{-x}+8e^{-2x}&=0\\
8e^{-2x}&=4e^{-x}\\
2e^{-2x}&=e^{-x}\\
2&=e^{-x}/e^{-2x}\\
2&=e^{x}\\
\ln(2) &=x
\end{aligned}"

To find the absolute maximum and absolute minimum, then, we evaluate "f" at the critical point and on the endpoints of the interval:

"\begin{aligned}
f(\ln(2))=&4e^{−\ln(2)}−4e^{−2\ln(2)}\\
=&4(\tfrac{1}{2})-4(\tfrac{1}{4})\\
=&2-1\\
=&1
\end{aligned}"

"\begin{aligned}
f(0)=&4e^{0}−4e^{0}\\
=&4-4\\
=&0
\end{aligned}"

"\begin{aligned}
f(1)=&4e^{−1}−4e^{−2}\\
=&4(0.3679)-4(0.1353)\\
=&1.4716-0.5412\\
\approx&0.93
\end{aligned}"

Therefore, "f" achieves its absolute minimum of at "x=0" and its absolute maximum of "1" at "x= \ln(2)"


Thus:

Absolute maxima = "1"

Absolute minima =